# Group Interventions on Deep Networks for Causal Discovery in Subsystems

### v1

@@ III-A Problem definition
We consider an $N$-variate time series, $Z^{N}=\{Z_{i}\}_{i=1}^{N}$, which are organized into $G$ distinct groups or subsystems, $X^{G}=\{X_{i}\}_{i=1}^{G}$, as illustrated in Figure 1. Each group $X_{i}$ is composed of $D_{i}$ variables. The total number of variables across all groups is the sum of the dimensions of the individual groups: $N=\sum_{i=1}^{G}D_{i}.$
@@ III-E Group-based Interventions
Here, $do(X_{i}=\tilde{X}_{i})$ represents an intervention where $X_{i}$ is replaced with $\tilde{X}_{i}$, a knockoff that is generated independent of the output. The term $E[X_{j}]$ captures the expected behavior of the target group under this intervention, isolating the causal influence of $X_{i}$ and facilitating the testing of causal relationships. The use of the $do$-operator [36] emphasizes that the intervention actively manipulates the input group of variables $X_{i}$ using their knockoff counterpart $\tilde{X}_{i}$, rather than passively observing their natural variations. The generation of knockoff variables $\tilde{X}_{i}$ follows the framework proposed in [4, 5, 6]. Knockoffs are constructed to serve as plausible but altered versions of the original variables.
@@ III-G Multi-set Canonical Correlation
where $\lambda$ represents the eigenvalues corresponding to the squared canonical correlations, and $A$ are the eigenvectors defining the optimal linear combinations. The eigenvectors corresponding to the largest eigenvalues are selected to maximize the inter-group correlations, ensuring the highest possible correlation between the canonical variables across groups. This aligns with the objective of CCA, which is to capture shared relationships between different groups of variables.
@@ IV-B Real-world Data
Data preparation is performed both at the individual site and network levels. For our experiments, we selected several measurement sites from the Fluxnet2015 dataset, including Hainich (DE-Hai), Monte Bondone (IT-MBo: Grasslands), Puechabon (FR-Pue: Evergreen Broadleaf Forests), and Tonzi Ranch (US-Ton: Woody Savannas). These sites were chosen for their ecological diversity, representing various ecosystem types across distinct geographic regions. The dataset includes climate and ecosystem time-series data such as global radiation ($R_{\text{g}}$), temperature ($T$), gross primary production ($GPP$), and ecosystem respiration ($R_{\text{eco}}$), recorded at multiple temporal resolutions (e.g., hourly, daily, weekly).
We categorize these variables into two groups: the climate group ($G_{\text{C}}$), comprising temperature ($T$) and global radiation ($R_{\text{g}}$), and the ecosystem group ($G_{\text{E}}$), consisting of the ecosystem variables $GPP$ and $R_{\text{eco}}$. We opt for daily sampling frequency, which offers advantages in mitigating the influence of daily patterns that might obscure the underlying causal relationships in the data.
@@ VI Conclusion
In this study, we introduced a deep learning-based approach to identify causal interactions within groups of variables or subsystems. By using deep networks for structure learning in an $N$-variate time series system, we perform invariance testing through group-based interventions to uncover causal relationships between groups of variables. Our method can identify bidirectional causal relationships, revealing mutual influences driven by feedback mechanisms in cause-effect systems. While our approach outperforms other causality methods on both synthetic and real-world datasets, it is important to note that the reliance on deep learning comes with a high computational cost. Our findings underscore the significance of data characteristics in causal discovery, emphasizing the need for adaptive methods since no single approach works universally. Nevertheless, gCDMI’s strong performance across various domains demonstrates its potential for broader applications in climate science, geophysics, and neuroscience. Future work will aim to improve the scalability of gCDMI and enhance its ability to handle non-stationarity through more generalizable models capable of adapting to regime shifts. Additionally, integrating domain expertise into the causal discovery process remains an important direction to ensure the interpretability and practical relevance of the results.

### v2

@@ III-A Problem definition
We consider an NN-variate time series, $Z^{N}=\{Z_{i}\}_{i=1}^{N}$, which are organized into $G$ distinct groups or subsystems, $X^{G}=\{X_{i}\}_{i=1}^{G}$, as illustrated in Figure 1. Each group $X_{i}$ is composed of $D_{i}$ variables. The total number of variables across all groups is the sum of the dimensions of the individual groups: $N=\sum_{i=1}^{G}D_{i}.$
@@ III-E Group-based Interventions
Here, $do(X_{i}=\tilde{X}_{i})$ represents an intervention where $X_{i}$ is replaced with $\tilde{X}_{i}$, a knockoff that is generated independent of the output. The term $E[X_{j}]$ captures the expected behavior of the target group under this intervention, isolating the causal influence of $X_{i}$ and facilitating the testing of causal relationships. The use of the dodo-operator [36] emphasizes that the intervention actively manipulates the input group of variables $X_{i}$ using their knockoff counterpart $\tilde{X}_{i}$, rather than passively observing their natural variations. The generation of knockoff variables $\tilde{X}_{i}$ follows the framework proposed in [4, 5, 6]. Knockoffs are constructed to serve as plausible but altered versions of the original variables.
@@ III-G Multi-set Canonical Correlation
where $\lambda$ represents the eigenvalues corresponding to the squared canonical correlations, and AA are the eigenvectors defining the optimal linear combinations. The eigenvectors corresponding to the largest eigenvalues are selected to maximize the inter-group correlations, ensuring the highest possible correlation between the canonical variables across groups. This aligns with the objective of CCA, which is to capture shared relationships between different groups of variables.
@@ IV-B Real-world Data
Data preparation is performed both at the individual site and network levels. For our experiments, we selected several measurement sites from the Fluxnet2015 dataset, including Hainich (DE-Hai), Monte Bondone (IT-MBo: Grasslands), Puechabon (FR-Pue: Evergreen Broadleaf Forests), and Tonzi Ranch (US-Ton: Woody Savannas). These sites were chosen for their ecological diversity, representing various ecosystem types across distinct geographic regions. The dataset includes climate and ecosystem time-series data such as global radiation ($R_{\text{g}}$), temperature ($T$), gross primary production (GPPGPP), and ecosystem respiration ($R_{\text{eco}}$), recorded at multiple temporal resolutions (e.g., hourly, daily, weekly).
We categorize these variables into two groups: the climate group ($G_{\text{C}}$), comprising temperature ($T$) and global radiation ($R_{\text{g}}$), and the ecosystem group ($G_{\text{E}}$), consisting of the ecosystem variables GPPGPP and $R_{\text{eco}}$. We opt for daily sampling frequency, which offers advantages in mitigating the influence of daily patterns that might obscure the underlying causal relationships in the data.
@@ VI Conclusion
In this study, we introduced a deep learning-based approach to identify causal interactions within groups of variables or subsystems. By using deep networks for structure learning in an NN-variate time series system, we perform invariance testing through group-based interventions to uncover causal relationships between groups of variables. Our method can identify bidirectional causal relationships, revealing mutual influences driven by feedback mechanisms in cause-effect systems. While our approach outperforms other causality methods on both synthetic and real-world datasets, it is important to note that the reliance on deep learning comes with a high computational cost. Our findings underscore the significance of data characteristics in causal discovery, emphasizing the need for adaptive methods since no single approach works universally. Nevertheless, gCDMI’s strong performance across various domains demonstrates its potential for broader applications in climate science, geophysics, and neuroscience. Future work will aim to improve the scalability of gCDMI and enhance its ability to handle non-stationarity through more generalizable models capable of adapting to regime shifts. Additionally, integrating domain expertise into the causal discovery process remains an important direction to ensure the interpretability and practical relevance of the results.